OpenStudy (anonymous):

find dy/dx if (y^2)-3xy+x^2+7

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

implicit diff yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

take the derivative with respect to x, remembering that y is a function of x. so for example the derivative of \[y^2\] is \[2yy'\] by the chain rule

OpenStudy (anonymous):

i assume it is set = 0 or something yes?

OpenStudy (anonymous):

\[y^2-3xy+x^2=7=0\]?

OpenStudy (anonymous):

i mean \[y^2-3xy+x^2+7=7\]

OpenStudy (anonymous):

damn

OpenStudy (anonymous):

=0 i mean not = 7

OpenStudy (anonymous):

you get \[2yy' -3y - 3xy'+2x=0\]

OpenStudy (anonymous):

then solve for y'

OpenStudy (anonymous):

\[2yy'-3xy'=3y-2x\] \[y'(2y-3x)=3y -2x\] \[y'=\frac{3y-2x}{2y-3x}\]